The following information relates to a joint production process for three products, with a total joint production cost of $155,000. There are no separable processing costs for any of the three products.

Product/ Sales Value of split off Units at Split-Off
1 $ 217,000 450    
2 46,500 1,050
3 46,500 1,500
$ 310,000 3,000

What percentage of joint cost is allocated to each of the three products using the sales value at split-off method?

P1: %, P2: %, P3: %

Subject: Cost management.
Answer #1

In case of sales value at split off method, joint costs are allocated on the basis of proportion of each products value to total sales value at split off. The number of physical units are irrelevant in this scenario. Since there is no processing, the given value is sales value at split off.

Sales Value of split off

% of total value

details

Joint cost allocation

Percentage allocated

P1

217,000

217,000/310,000 = 70%

155000 * 70%

                         108,500

108,500/155,000 = 70%

P2

46,500

46500/310000 = 15%

155000 * 15%

                            23,250

23,250/155,000 = 15%

P3

46,500

15%

155000 * 15%

                            23,250

23,250/155,000 = 15%

310,000

                         155,000

100%

Alternatively, the percentage of sales value at split off is the percentage at which joint costs allocated i.e. 70% , 15% and 15%. The above table simplifies the understanding.

Kindly upvote